0.15mg. the drug comes in 0.05 mg per ml. Give ml

125 mg per 5 ml. The order is for 0.25 Gm. Give in ml.

500 mg. per 5 cc. 250 mg are ordered. Give in ml

6,000 units subcutaneously three times a week. The label reads Epogen 20,000 units/ml. How many millilitres should the nurse give the patient?

0.15 mg. The drug comes in 0.05 mg per ml. Give in ml.

125 mg per 5ml. The order is for 0.25Gm. Give in ml.
Answer #1

ANSWER.

* Calculation of ml of drug required.

       Formula = Desired dose /Dose in hand * Quantity.

   Here

      Desired dose (D) = 0.15 mg

      Dose in hand (H) = 0.05 mg

          Quantity (Q) = 1 ml

As per above formula

    Volume(ml) of drug required = D/H * Q

                                                  = 0.15/0.05 * 1

                                                  = 3 * 1

                                                  = 3 ml

* Calculation of ml of drug required.

      Formula = Desired dose /Dose in hand * Quantity.

Here

      Desired dose (D) = 0.25 gm.

      Dose in hand (H) = 125 gm.

          Quantity (Q) = 5 ml

As per above formula

     Volume(ml) of drug required = D/H * Q

                                                   = 0.25/125 * 5

                                                   = 0.002 * 5

                                                   = 0.01 ml

* Calculation of ml of drug required.

    Formula = Desired dose /Dose in hand * Quantity.

Here

        Desired dose (D) = 250 mg

        Dose in hand (H) = 500 mg

            Quantity (Q) = 5 ml

As per above formula

       Volume(ml) of drug required = D/H * Q

                                                     = 250/500 * 5

                                                     = 0.5 * 5

                                                     = 2.5 ml.

* Calculation of ml of drug required.

        Formula = Desired dose/Doe in hand * Quantity.

   Here

        Desired dose (D) = 6000 units.

        Dose in hand (H) = 20000 units.

            Quantity (Q) = 1 ml

As per above formula

        Volume(ml) of drug required = D/H * Q

                                                      = 6000/20000 * 1

                                                      = 0.3 * 1

                                                      = 0.3 ml